### v3

@@ Abstract
Street View Imagery (SVI) is a valuable data source for studies (e.g., environmental assessments, green space identification or land cover classification). While commercial SVI is available, such providers commonly restrict copying or reuse in ways necessary for research. Open SVI datasets are readily available from less restrictive sources, such as Mapillary, but due to the heterogeneity of the images, these require substantial preprocessing, filtering, and careful quality checks. We present an efficient method for automated downloading, processing, cropping, and filtering open SVI, to be used in a survey of human perceptions of the streets portrayed in these images. We demonstrate our open-source reusable SVI preparation and smartphone-friendly perception-survey software with Amsterdam (Netherlands) as the case study. Using a citizen science approach, we collected from 331 people 22,637 ratings about their perceptions for various criteria. We have published our software in a public repository for future re-use and reproducibility.
@@
Figure 1: Sample image from Amsterdam (source: Mapillary)
@@ 1 Introduction
Nonetheless, most urban design, transportation planning and environmental epidemiology researchers rely on objectively measured streetscape indicators because it is easier to collect measurable and quantifiable physical attributes, such as for indicators of green spaces (Kothencz and Blaschke, 2017; Labib et al., 2020; Liu and Sevtsuk, 2024). However, people’s subjective feelings about places do not necessarily correlate with such objectively measured spatial indicators (McCrea et al., 2006). To capture those sentiments as usable data, researchers typically use resource-intensive methods such as field interviews or questionnaires (Lynch, 1960; Lotfi and Koohsari, 2009). Therefore, we argue that developing an efficient yet easy-to-implement method to capture human perceptions of urban environments would provide a much needed tool for answering questions like those posed above. Taken further, these answers can help with efforts to achieve the United Nations Sustainable Development Goals111www.un.org/sustainabledevelopment/sustainable-development-goals/ such as Goal 11, ‘Sustainable Cities and Communities’.
Over the past decade, various companies, organisations and people have been collecting Street View Imagery (SVI). Since the launch of Google Street View in 2007 (Vincent, 2007), the resulting datasets of SVI have been successfully used in numerous studies over a wide variety of domains (Biljecki and Ito, 2021). Some of the earliest research validating the usefulness of virtual streetscape audits in place of physical audits (Badland et al., 2010). Looking back at a collection of more recent work over the past decade, Dai et al. (2024, pp. 16–17) found that SVI ‘performs exceptionally well in capturing environmental variables’ and ‘holds immense potential to facilitate environmental health-related studies in the big data era’. However, we find challenges remain in the use of SVI for research purposes. In particular, we focus on two major classes of SVI currently accessible to researchers: (i) commercial SVI such as offered by Google, Bing, Baidu or Tencent and (ii) Volunteered Street View Imagery (VSVI) from open platforms such as Mapillary and KartaView (formerly OpenStreetCam), which collect geo-tagged volunteer-submitted photographs and make them available from a central repository.
Unlike commercial SVI platforms like Google, which publish their content under licenses that prohibit certain research methods, VSVI platforms such as Mapillary permit usage of their imagery under the Creative Commons Attribution-ShareAlike 4.0 International222creativecommons.org/licenses/by-sa/4.0/deed.en license, making it easy to adapt into an open-source software system for collecting human perception responses. However, the quality of VSVI can be vary considerably from very poor to excellent (Ma et al., 2019). Recent work by Zheng and Amemiya (2023) has highlighted ways to improve the usability of VSVI by undergoing a process of filtering, however, no studies have yet developed a comprehensive open-source toolkit to use such filtered VSVI in human perception research.
@@ 2 Background
All of these aforementioned works relied on either author-supplied photographs or commercial SVI, usually Google, but some used Tencent or Baidu. We instead set out to build reusable, open and FAIR software for assembling and operating street view perception surveys based on an open provider of VSVI, Mapillary.
@@ 3.1 Mapillary street view imagery
Our method uses imagery from the Mapillary platform, which offers free-to-use SVI and VSVI in many cities around the world under the terms of a Creative Commons license. Some of this imagery was collected professionally with high-quality 360-degree panoramic cameras, especially in cities like Amsterdam (see Figure 1 for a sample of a cropped subimage of a panoramic photograph). Most of the available imagery is VSVI, of varying quality, but recent work by Zheng and Amemiya (2023) shows that such VSVI can be usefully filtered and employed in research as an alternative to higher-cost or non-free options.
The motivation to use Mapillary VSVI comes from the FAIR data movement (Wilkinson et al., 2016). Guidance for the final FAIR principle, reusability, emphasizes that data must be not only technically interoperable but also legally interoperable333www.go-fair.org/fair-principles/, section R1.1. However, the standard terms of service444cloud.google.com/maps-platform/terms/ for Google SVI stresses that images are non-free and explicitly forbid the kind of usage that we need for research. In particular, in section 3.2.3 of the terms of service, they prohibit scraping, pre-fetching, bulk downloading, storing, and resharing Google SVI, all of which are necessary components of our survey method. Bing, Baidu and Tencent Maps also list several similar restrictions in their terms of service, and in the case of the latter two, those platforms are mainly focused on providing imagery from China rather than worldwide.
While Google imagery is high-quality and widely available, we intend to keep within the open and FAIR principles. In particular, we are concerned that some people wishing to use our toolkit may not be able to obtain the authorization to use Google imagery, or to produce freely-reusable datasets from it, such as those who work in countries where there is legal uncertainty about these ways of using proprietary data (Helbich et al., 2024). We simply wish to show that there are alternatives that are free in the sense of being accessible and available (‘libre’) for research usage without significant restrictions on reusability. The fact that Mapillary is free in the sense of monetary cost (‘gratis’) is also helpful. In contrast to Google, the Mapillary terms of service555www.mapillary.com/terms allows usage of VSVI under a Creative Commons license that only require proper attribution of imagery that is downloaded and reshared or integrated into applications.
@@ 3.2 Imagery downloading and processing
Our Python script, mapillary_jpg_download.py gathers imagery from Mapillary via their API (see Appendix C.1 for more details). The script only requires that you obtain your own free API key from the Mapillary developer portal, and that you provide a geographic bounding rectangle for your region of interest. It is also possible to obtain imagery from other sources, and inject it into the process at this point, should Mapillary be unsuitable or undesired for any reason, however the focus of this work is on Mapillary-provided imagery.
Once the tile data files and photographic imagery are downloaded, we must process a very large number of JPEG files, some of which are panoramic photographs projected into a wide image format, and others which are simply plain photographs. The two main tasks we need to accomplish are: (1) finding ‘sensible subimages’ in the panoramic imagery to crop out and save separately, and (2) weeding out any images that are too dark, too blurry or defective in some other way (such as only showing a wall or an undifferentiated block of greenery). These are both somewhat vague needs but thanks to recent advances in machine learning they can both be addressed using off-the-shelf software. In the case of panoramic photographs, we took subimages of a 4:3 ratio to show users images that fit within the app’s available view-port for imagery, rather than distorted and overly large raw panoramic images.
@@ 3.3 Filtering by image quality
Most of the volunteered photographs that we encountered on Mapillary are non-panoramic and can range from very good quality to quite poor. Inspired by the filtering ideas of Zheng and Amemiya (2023), we calculate contrast and ‘tone-mapping score’ (Stefanescu, 2021) using the Python Scikit-image library (Van der Walt et al., 2014), and apply a certain threshold on a formula composing both values (see Appendix C.5), to select acceptable images. We also applied the road center-line finding algorithm described above, but in this case, only to determine if there is a road or not within the image, and whether that road center-line lies closer to the center of the image than to the edge. Images that pass these tests are then cropped to fit a 4:3 ratio so they fit into the image view-port of our web app.
@@ 3.4 Survey frontend
The perception-gathering survey is written as a mobile web app in the ReactJS framework. First of all, before the main perception-gathering survey takes place, there is a short socio-demographic survey requesting a few personal details: age, education level, gender, approximate income, home postal code, country of residence and consent for data collection (see Section A.2 for more details). Only age and consent are required fields. We comply with the EU General Data Protection Regulations666commission.europa.eu/law/law-topic/data-protection_en and require participants’ explicit consent, which may be withdrawn at any time. Furthermore, we do not use data from participants who are under 18.
The main perception-gathering survey then proceeds: it shows participants one image at a time along with one of five possible categories: walkability, bikeability, pleasantness, greenness or safety (see Appendix A.1 for more details). Participants may rate each image according to the given category by pressing one of five rating buttons, or swiping the image towards one of the rating buttons. The buttons are arranged along a Likert-type scale: awful, bad, neutral, good, and great (internally numbered from 1 to 5). Participants rate five images according to a given category and then the app chooses a new category at random, until 20 images have been rated in each category. There are also options to skip the current image, undo the last rating (see Appendix B.2 for more undo details), go full screen, change language, get more help, and report an issue. The mobile web app (Figure 4) is designed to fit seamlessly on a mobile device so that users can pull up their smartphone to quickly do some ratings whenever they desire.
The frontend runs entirely on the user’s browser and communicates with the backend server via a public API (see Appendix B.1 for more API details). Image URLs are fetched from the backend, and then downloaded from the image-hosting server as needed. Ratings and undo commands are sent asynchronously to the backend while the interface updates.
@@ 3.5 Survey backend
The backend is an independent ExpressJS-based server with a well-defined public API for tasks such as starting new sessions and submitting surveys and ratings. The backend is a separate module from the frontend; it is possible that a different frontend (or even raw requests) could interface with the backend, if so desired. The public API (see Section B.1) is treated as a potential entry point for input from any possible source, including malicious ones. Data is stored securely on a private server using PostgreSQL and PostGIS for geographic data processing. The overall arrangement of frontend, backend and client (in this case, depicted as a smartphone user) is shown in Figure 5.
@@ 4.2 Collected perception
estimated monthly gross income
@@
Of our 331 participants, 59.2% (196) submitted at least 50 ratings, and 50.8% (168) finished the full 100. The median survey completion time was 6.3 minutes and 93.4% (309) of participants took 30 minutes or less from start to end. We observed participants swiping or clicking with ease through our smartphone-friendly user interface, and we later confirmed this using timestamps in our database: the median interval between ratings was 3.7 seconds and 89.7% (20,310) of the submitted ratings occurred within 10 seconds of the previous rating.

### v4

@@ Abstract
Street View Imagery (SVI) is a valuable data source for studies (e.g., environmental assessments, green space identification or land cover classification). While commercial SVI is available, such providers commonly restrict copying or reuse in ways necessary for research. Open SVI datasets are readily available from less restrictive sources, such as Mapillary, but due to the heterogeneity of the images, these require substantial preprocessing, filtering, and careful quality checks. We present a method for automated downloading, processing, cropping, and filtering open SVI, to be used in a survey of human perceptions of the streets portrayed in these images. We demonstrate our open-source reusable SVI preparation and smartphone-friendly perception-survey software with Amsterdam (Netherlands) as the case study. Using a citizen science approach, we collected from 331 people 22,637 ratings about their perceptions for various criteria. We have published our software in a public repository for future re-use and reproducibility.
@@ 1 Introduction
Nonetheless, most urban design, transportation planning and environmental epidemiology researchers rely on objectively measured streetscape indicators because it is easier to collect measurable and quantifiable physical attributes, such as for indicators of green spaces (Kothencz and Blaschke, 2017; Labib et al., 2020; Liu and Sevtsuk, 2024). However, people’s subjective feelings about places do not necessarily correlate with such objectively measured spatial indicators (McCrea et al., 2006). To capture those sentiments as usable data, researchers typically use resource-intensive methods such as field interviews or questionnaires (Lynch, 1960; Lotfi and Koohsari, 2009). Therefore, we argue that developing a readily deployable method to capture human perceptions of urban environments would provide a much needed tool for answering questions like those posed above. Taken further, these answers can help with efforts to achieve the United Nations Sustainable Development Goals111www.un.org/sustainabledevelopment/sustainable-development-goals/ such as Goal 11, ‘Sustainable Cities and Communities’.
Over the past decade, various companies, organisations and people have been collecting Street View Imagery (SVI) since Google began in 2007 (Vincent, 2007) and the resulting datasets have been widely studied (Biljecki and Ito, 2021). However, even after all this time, we find challenges remain in the use of SVI for research purposes. In particular, we focus on two major classes of SVI currently accessible to researchers: (i) commercial SVI such as offered by Google, Bing, Baidu or Tencent and (ii) Volunteered Street View Imagery (VSVI) from open platforms such as Mapillary and KartaView (formerly OpenStreetCam), which collect geo-tagged volunteer-submitted photographs and make them available from a central repository.
Unlike commercial SVI platforms like Google, which publish their content under licenses that prohibit certain research methods, VSVI platforms such as Mapillary permit usage of their imagery under the Creative Commons Attribution-ShareAlike 4.0 International222creativecommons.org/licenses/by-sa/4.0/deed.en license, making it easy to adapt into an open-source software system for collecting human perception responses. A common complaint is that the quality of VSVI can vary considerably from very poor to excellent (Ma et al., 2019), but this problem can be ameliorated by processing and filtering (Zheng and Amemiya, 2023). However, no studies have yet developed an open-source toolkit that downloads, filters and processes VSVI and configures a mobile (web) app-based survey for human perception research using it.
@@ 2 Background
Since the launch of Google Street View in 2007 (Vincent, 2007), the resulting datasets of SVI have been effectively used in numerous studies over a wide variety of domains from urban planning to public health (Biljecki and Ito, 2021). Some of the earliest research validating the usefulness of virtual streetscape audits in place of physical audits (Badland et al., 2010). Looking back at a collection of more recent work over the past decade, Dai et al. (2024, pp. 16–17) found that SVI ‘performs exceptionally well in capturing environmental variables’ and ‘holds immense potential to facilitate environmental health-related studies in the big data era’. Mapillary is a major source of VSVI and includes high-quality panoramic images. With those there is the possibility to crop subimages based on street morphology (Beaucamp et al., 2022). However, Mapillary also contains many non-panoramic images and the quality of those can be highly variable (Hou and Biljecki, 2022; Ma et al., 2019). Recent work by Zheng and Amemiya (2023) has highlighted ways to improve the usability of VSVI by undergoing a process of filtering, while Biljecki et al. (2023) found that the use of non-panoramic images only slightly detracted from their results for greenness and sky view visibility indices in the cases they studied compared to the same tests run on panoramic images. Hou and Biljecki (2022) developed a framework for defining SVI and a consistent method to evaluate its quality. Ding et al. (2021) used Mapillary image sequences to find bikeway networks in Malmö using sign detection machine learning techniques and Sánchez and Labib (2024) has presented a toolkit using Mapillary images for measuring greenness visibility at eye level.
@@
The user interface presented by our survey app utilizes a game-like ‘design pattern’ (Morschheuser et al., 2017) with large icons that can be swiped, and congratulatory messages at various stages of progress, but does not proceed any further down that gamification path. It is not as elaborate as MapSwipe (Ullah et al., 2023) or Tomnod (Baruch et al., 2016), both of which involve crowdsourcing spatial data that is rendered on a map, but rather more akin to Galaxy Zoo (Masters et al., 2019), in which images are categorized by clicking a button (in that case, by galaxy type) in a slick and customized (web) app.
All of these aforementioned SVI-based works relied on either author-supplied photographs or commercial SVI, usually Google, but some used Tencent or Baidu. We instead set out to build reusable, open and FAIR software for assembling and operating street view perception surveys based on an open provider of VSVI, Mapillary.
@@ 3.1 Mapillary street view imagery
Figure 1: Sample image from Amsterdam (source: Mapillary)
Our method uses imagery from the Mapillary platform, which offers free-to-use SVI and VSVI in many cities around the world under the terms of a Creative Commons license. Some of this imagery was collected professionally with high-quality 360-degree panoramic cameras, especially in cities like Amsterdam (see Figure 1 for a sample of a cropped subimage of a panoramic photograph). Most of the available imagery is VSVI, of varying quality, but recent work by Zheng and Amemiya (2023) shows that such VSVI can be usefully filtered and employed in research as an alternative to higher-cost or non-free options, while Biljecki et al. (2023) found that working with non-panoramic imagery could bring them reasonably close to deriving the same results as working with panoramic imagery in the cases they studied. All of Mapillary’s available imagery is processed with face- and license plate-blurring software according to their privacy policy333www.mapillary.com/privacy. However, whichever SVI provider a researcher chooses to use, they should ensure that the images are appropriate, ethically sound and privacy protected, as they would with any other data source.
The motivation to use Mapillary VSVI comes from the FAIR data movement (Wilkinson et al., 2016). Guidance for the final FAIR principle, reusability, emphasizes that data must be not only technically interoperable but also legally interoperable444www.go-fair.org/fair-principles/, section R1.1. However, the standard terms of service555cloud.google.com/maps-platform/terms/ for Google SVI stresses that images are non-free and explicitly forbid the kind of usage that we need for research. In particular, in section 3.2.3 of the terms of service, they prohibit scraping, pre-fetching, bulk downloading, storing, and resharing Google SVI, all of which are necessary components of our survey method. Bing, Baidu and Tencent Maps also list several similar restrictions in their terms of service, and in the case of the latter two, those platforms are mainly focused on providing imagery from China rather than worldwide.
While Google imagery is high-quality and widely available, we intend to keep within the open and FAIR principles. In particular, we are concerned that some people wishing to use our toolkit may not be able to obtain the authorization to use Google imagery, or to produce freely-reusable datasets from it, such as those who work in countries where there is legal uncertainty about these ways of using proprietary data (Helbich et al., 2024). We simply wish to show that there are alternatives that are free in the sense of being accessible and available (‘libre’) for research usage without significant restrictions on reusability. The fact that Mapillary is free in the sense of monetary cost (‘gratis’) is also helpful. In contrast to Google, the Mapillary terms of service666www.mapillary.com/terms allows usage of VSVI under a Creative Commons license that only require proper attribution of imagery that is downloaded and reshared or integrated into applications. We also note that there are some countries for which Google SVI is sparse or barely available at all, while Mapillary has substantially more VSVI in those same countries that can be reasonably filtered and processed; e.g., as of September 2024 this is the case in Zambia, Morocco and Nicaragua, among others.
@@ 3.2 Imagery downloading and processing
Our Python script, mapillary_jpg_download.py gathers imagery from Mapillary via their API (see Appendix C.1 for more details). The script only requires the acquisition of a free API key from the Mapillary developer portal, and to be given a geographic bounding rectangle for some region of interest. It is also possible to obtain imagery from other sources, and inject it into the process at this point, should Mapillary be unsuitable or undesired for any reason, however the focus of this work is on Mapillary-provided imagery.
Once the tile data files and photographic imagery are downloaded, we must process a very large number of JPEG files, some of which are panoramic photographs projected into a wide image format, and others which are simply plain photographs. The two main tasks we need to accomplish are: (1) finding ‘sensible subimages’ in the panoramic imagery to crop out and save separately, and (2) weeding out any images that are too dark, too blurry or defective in some other way (such as only showing a wall or an undifferentiated block of greenery). These are both somewhat vague needs but thanks to recent advances in machine learning they can both be addressed using off-the-shelf software. In the case of panoramic photographs, our toolkit automatically crops subimages of a 4:3 ratio to show users images that fit within the app’s available view-port for imagery, rather than distorted and overly large raw panoramic images.
@@ 3.3 Filtering by image quality
Most of the volunteered photographs that we encountered on Mapillary, while looking at various cities around the world, are non-panoramic and can range from very good quality to quite poor. For example, some unusable images we found barely showed anything at all but a single wall, a blur of greenery or a close-in view of a parked car. Inspired by the filtering ideas of Zheng and Amemiya (2023), we calculate contrast and ‘tone-mapping score’ (Stefanescu, 2021) using the Python Scikit-image library (Van der Walt et al., 2014), and apply a certain threshold on a formula composing both values (see Appendix C.5), to select acceptable images. We also applied the road center-line finding algorithm described above, but in this case, only to determine if there is a road or not within the image; through trial and error, we found this simple ‘road check’ to be a good heuristic that met our need to capture an open and clear view of a street and the surroundings that a person would see if they were standing there. Images that pass these tests are then automatically cropped by our toolkit to a 4:3 ratio so they fit into the image view-port of our web app. However, all of these the filtering criteria can be adjusted, or modified entirely, to suit any particular research need. For example, if a researcher would like to include ‘non-road’ images in their dataset, then that is possible by turning off the road check.
@@ 3.4 Survey frontend
The perception-gathering survey is written as a mobile web app in the ReactJS framework. First of all, before the main perception-gathering survey takes place, there is a short socio-demographic survey requesting a few personal details: age, education level, gender, approximate income, home postal code, country of residence and consent for data collection (see Section A.2 for more details). Only age and consent are required fields. We comply with the EU General Data Protection Regulations777commission.europa.eu/law/law-topic/data-protection_en and require participants’ explicit consent, which may be withdrawn at any time. Furthermore, we do not use data from participants who are under 18.
The main perception-gathering survey then proceeds: it shows participants one image at a time along with one of five possible categories: walkability, bikeability, pleasantness, greenness or safety (see Appendix A.1 for more details). Participants may rate each image according to the given category by pressing one of five rating buttons, or swiping the image towards one of the rating buttons. The buttons are arranged along a Likert-type scale: awful, bad, neutral, good, and great (internally numbered from 1 to 5). Participants rate five images according to a given category and then the app chooses a new category at random, until 20 images have been rated in each category. There are also options to skip the current image, undo the last rating (see Appendix B.2 for more undo details), go full screen, change language, get more help, and report an issue. The mobile web app (Figure 4) is designed to fit seamlessly on a mobile device so that users can pull up their smartphone to quickly do some ratings whenever they desire. Our decision to collect absolute scores (instead of pairwise comparisons) makes it possible for us to present a clean and simple user interface to participants, with an intuitive swipe-to-rate input mechanism.
The frontend runs entirely on the user’s browser and communicates with the backend server via a public API (see Appendix B.1 for more API details). Image URLs are fetched from the backend, and then downloaded from the image-hosting server as needed. Ratings and undo commands are sent asynchronously to the backend while the interface updates. Text in the user interface is managed by the react-intl-universal module so that we (or other contributors) can easily add new language translations and locales when preparing to deploy the survey in different countries.
@@ 3.5 Survey backend
The backend is an independent ExpressJS-based server with a well-defined public API for tasks such as starting new sessions and submitting surveys and ratings. The backend is a separate module from the frontend; it is possible that a different frontend (or even raw requests) could interface with the backend, if so desired. The public API (see Section B.1) is treated as a potential entry point for input from any possible source, including malicious ones. Data are stored securely on a private server using PostgreSQL and PostGIS for geographic data processing, where they are kept until they have been suitably anonymized or stripped of personally-identifiable information in a manner that is appropriate for publication. The overall arrangement of frontend, backend and client (in this case, depicted as a smartphone user) is shown in Figure 5.
@@ 4.2 Collected perception
estimated monthly gross income (mgi)
@@
id timestamp sess image cat score postcode country age mgi education gender
8710 2023-06-27 16:21:03 145 77114 3 4 3--- -- Netherlands 26 2300 Postgraduate man
8719 2023-06-27 16:21:23 151 85673 2 3 3--- Netherlands 38 4000 Postgraduate woman
8976 2023-06-27 16:43:48 151 71732 1 4 3--- Netherlands 38 4000 Postgraduate woman
13378 2023-07-12 15:05:25 211 42535 4 2 1--- -- Nederland 35 3500 Tertiary man
13460 2023-07-12 22:47:45 212 48075 4 2 1--- -- netyerlands 58 5000 Postgraduate woman
19100 2023-09-12 10:33:38 301 190126 4 4 1--- Belgium 18 200 Secondary non-binary
where
id: A unique identifier assigned to every rating submitted by any participant. timestamp: The date and time of the submission (recorded to the microsecond, not shown here) sess: The unique identifier assigned to every participant when they start a rating session. image: The unique identifer (previously configured in our database) of the image that was rated. cat: The rating category: 1=Walkability, 2=Bikeability, 3=Pleasantness, 4=Greenness, 5=Safety. score: The submitted rating: 1=awful, 2=bad, 3=neutral, 4=good, 5=great. postcode: Home postal code of the participant (redacted here). Free-form entry in the survey. country: Country of the participant. Free-form entry in the survey, hence we must deal with variation. age: Age of the participant. Aside from consent, the only required entry in the survey, and numeric. mgi: Estimated monthly gross income. Optional in our survey, localized to the Netherlands (€). education: One of: Primary, Secondary, Tertiary or Postgraduate. Optional in our survey. gender: One of: woman, non-binary, man, unspecified or free-form text entry.
Table 2: Data sample. Note that the underlying data is stored in normalized tables, this is a joined-together view for consideration and analysis. The postal codes have been mostly redacted, but their shape remains (with dashes replacing digits and letters) to show the variability of the underlying survey data. The timestamps and ages have been randomly perturbed as well (Rahman et al., 2023). As shown, the raw data can be messy and require some processing to clean up inconsistencies like those found above in the country column.
@@
Of our 331 participants, 59.2% (196) submitted at least 50 ratings, and 50.8% (168) finished the full 100. The median survey completion time was 6.3 minutes and 93.4% (309) of participants took 30 minutes or less from start to end. We observed participants swiping or clicking with ease through our smartphone-friendly user interface, and we later confirmed this using timestamps in our database: the median interval between ratings was 3.7 seconds and 89.7% (20,310) of the submitted ratings occurred within 10 seconds of the previous rating. A sample of the collected information is shown in Table 2.